You have a container of ice at 0C. What is the minimum mass of ice that must be added to 150.0g of liquid water at 15.0C to lower its temperature to 0.0C?
1
Expert's answer
2012-12-10T10:56:56-0500
Melting ice take the heat from water, hence
L*m(ice) = c*t*m(water)
where L = 335 kJ/kg is specific heat of melting for ice andc = 4200 J/(kg*K) is specific heat capacity.
m(ice) = c*t*m(water)/L = 4200*15*0.15/(335*10^(3)) = 0.0282kg = 28.2 g